# BConformeR: A Conformer Based on Mutual Sampling for Unified Prediction of Continuous and Discontinuous Antibody Binding Sites

### v1

@@ Abstract
Accurate prediction of antibody-binding sites (epitopes) on antigens is crucial for vaccine design, immunodiagnostics, therapeutic antibody development, antibody engineering, research into autoimmune and allergic diseases, and for advancing our understanding of immune responses. Despite in silico methods that have been proposed to predict both linear (continuous) and conformational (discontinuous) epitopes, they consistently underperform in predicting conformational epitopes. In this work, we propose a conformer-based model trained on antigen sequences derived from 1,080 antigen-antibody complexes, leveraging convolutional neural networks (CNNs) to extract local features and Transformers to capture long-range dependencies within antigen sequences. Ablation studies demonstrate that CNN enhances the prediction of linear epitopes, and the Transformer module improves the prediction of conformational epitopes. Experimental results show that our model outperforms existing baselines in terms of PCC, ROC-AUC, PR-AUC, and F1 scores on conformational epitopes.
Keywords B-cell Epitope Prediction; Conformer; Antigen–Antibody Interface
@@ 1 Introduction
Figure 1: Overview of epitope classification and BConformerR prediction. (A) Epitopes, i.e., Antibody-binding sites, include conformational fragments (n=2,632) and linear regions (n=2,388). Data derived from 1,432 antigen records. (B) Epitope prediction workflow: Antigen inputs are processed to predict linear and conformational epitopes, which enable applications in vaccine design and diagnostic development.
@@
Figure 2: BConformeR overview. (A) Preparation pipeline of the data. 1300 antigen-antibody FASTA files were processed using AlphaFold Multimer v3 to predict structures, with top-ranked structure for each was selected to extract epitopes (right); Antigen sequences were embedded by ESM-2, producing embeddings of dimension 1024×1280 (left). (B) Conformer architecture. Stem module applies convolution and max-pooling to the original ESM-2 embeddings; The Conv block and Trans block are based on ResNet bottleneck and Vision Transformer modules respectively; Feature coupling units (FCU-Up and FCU-Down) connect the Conv and Trans blocks, with 11 FCU layers in total; The classifier linearly fuses features from both Conv and Trans branches.
@@
Here, we propose BConformeR, a Conformer-based model for linear and conformational BCR epitope prediction. Antigen sequences from 1,180 Ag-Ab complexes were embedded using ESM-2, with 1,080 used for training and 100 for evaluation. In a blind test of 24 antigen records, BConformeR outperforms existing baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including PCC, ROC-AUC, PR-AUC, and F1 score on conformational epitopes. To further dissect the model’s design, we ablated ResNet bottlenecks and ViT modules from Conformer, showing that convolutional layers improve linear epitope prediction, and Transformer modules are critical for capturing discontinuous epitope patterns.
@@ 2.1 Antibody-agnostic prediction models
Numerous models, trained on antigen–epitope pairs, have been proposed and refined to enhance epitope prediction. SEPPA-3.0 was the epitope prediction tool specifically designed for glycoproteins, based on the logistic regression classifier SEPPA3.0 . BepiPred-3.0 used embeddings from pretrained protein language models ESM-2 as input to a feedforward neural network (FFNN) to predict both linear and conformational B-cell epitopes BepiPred3.0 . Epitope3D modeled antigen surfaces as residue-level graphs and extracted distance-based graph signatures to train classifiers for conformational B-cell epitope prediction, with AdaBoost selected as the backbone Epitope3D ; AdaBoost . ScanNet was an end-to-end geometric deep learning model that learned interpretable residue-level features directly from 3D structural neighborhoods ScanNet . SEMA-1D 2.0 was constructed by adding a fully-connected linear layer on an ensemble of five ESM-2 models, while SEMA-3D 2.0 was developed using an ensemble of five pre-trained Structure-aware Protein language models (SaProt) SEMA2.0 ; SaProt . DiscoTope-3.0 utilized inverse folding representations from ESM-IF1 and was trained on both predicted and solved structures using a positive-unlabelled ensemble strategy, enabling structure-based B-cell epitope prediction at scale DiscoTope3.0 ; ESM-IF1 .
@@ 3.1.2 Epitope classification
Generally, epitopes are separated into multiple Regions (R) and Fragments (F) for each antigen sequence. Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues. Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl . Fragments, which do not contain enough antibody-contacting residues to qualify as a region, are considered discontinuous in our study (Fig. 1A).
@@
Table 1: Architecture of BConformeR, a 1D adaptation of Conformer Conformer21 . The MHSA-12 indicates multi-head self-attention with 12 heads operating on 1536-dimensional embeddings (128 dimensions per head). $\mathrm{Conv}_{\mathrm{cls}}$ denotes the 1D convolutional classifier operating on CNN features, and $\mathrm{Trans}_{\mathrm{cls}}$ denotes the linear classifier operating on Transformer outputs. The final prediction is obtained by a weighted fusion of the two.
@@ 3.2.2 CNN branch
The channel dimension is first downsampled to half, then expanded by a factor of 4 within each bottleneck. The first convolution block contains a single bottleneck, while subsequent blocks include two bottlenecks: one applied before the Transformer branch, and the other afterward. Since convolution kernels slide over representation sequences with overlap, the CNN blocks are well-suited to capture residue-level contextual information, providing local features for the Transformer branch.
@@ 3.2.3 Transformer branch
where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 12) heads, and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$. The $\mathrm{DropPath}$ operator implements stochastic depth regularization.
@@ 3.3.1 AgIoU
To evaluate the overlap between the prediction and real epitopes, we use the Antigen-level Intersection over Union (AgIoU), which is defined as the ratio of the number of correctly predicted epitope residues to the total number of residues that are either labeled as epitope in the ground truth or prediction as such:
@@ 4.1.1 Dataset
We collected 1,300 Ab–Ag complex FASTA files with AlphaFold-2.3 colabfold version from AbEpiTope-1.0. Each complex was then modeled with AlphaFold2 Multimer v3, generating 30 structural predictions per complex and a total of 39,000 structures. For each complex, we selected the top-ranked model and excluded those with a pLDDT score below 0.82. Interface residues were extracted based on a distance cutoff of 4 Å between heavy atoms on the antigen and antibody chains (1). To enhance data quality and limit parameter overhead, we removed complexes containing insufficient epitope residues or antigens exceeding 1024 residues in length. Finally, the dataset contains 1,180 Ag-Ab complexes in total: 1,080 used for training BConformeR, and 100 set aside for evaluation and ablation. And the antigen chains annotated with epitope labels serve as the input.
@@
SEPPA-3.0 0.100 ± 0.020 0.606 ± 0.055 0.024 ± 0.001 0.058 ± 0.014 0.067 0.088 ± 0.005
BepiPred-3.0 0.210 ± 0.020 0.657 ± 0.032 0.034 ± 0.003 0.191 ± 0.003 0.215 0.149 ± 0.003
SEMA-1D 2.0 0.510 ± 0.030 0.683 ± 0.027 0.029 ± 0.001 0.182 ± 0.002 0.217 0.143 ± 0.002
DiscoTope-3.0 0.300 ± 0.020 0.706 ± 0.015 0.047 ± 0.003 0.227 ± 0.007 0.276 0.167 ± 0.002
BConformeR 0.300 ± 0.020 0.673 ± 0.014 0.105 ± 0.004 0.302 ± 0.003 0.347 0.212 ± 0.002
Table 2: Performance comparison of different methods on a blind test of 24 antigen records. F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively; Thresholds were selected following those of baseline methods; Standard deviation calculated over a range of thresholds.
Due to the differences in training and test datasets across existing epitope prediction methods, the selection of a blind test set was important for fair evaluation. Therefore, we also extracted 24 antigen records from the test sets of SEMA-1D 2.0 and DiscoTope-3.0. At least 75% of these records were previously used as test data by each baseline method, and 100% blind for BConformeR.
@@
Conformer-12 0.965 0.516 0.672 0.064 0.393 0.110 302.24 309.93
Conformer-9 0.946 0.489 0.645(-4.0%) 0.055 0.344 0.095(-13.6%) 230.27 236.12
Conformer-6 0.977 0.449 0.615(-8.5%) 0.041 0.263 0.071(-35.5%) 158.30 162.31
ResNet-152 0.979 0.514 0.674(+0.3%) 0.046 0.290 0.079(-28.2%) 242.44 248.62
ResNet-101 0.961 0.468 0.629(-6.4%) 0.027 0.216 0.048(-56.4%) 179.24 183.80
ViT-12 0.969 0.477 0.639(-4.9%) 0.049 0.355 0.086(-21.8%) 230.24 234.42
ViT-9 0.958 0.455 0.617(-8.2%) 0.047 0.328 0.082(-25.5%) 173.57 176.32
Table 3: Ablation study of different models evaluated on 100 antigen records. Conformer-12 corresponds to BConformeR, and Conformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to Conformer-12/9. ResNet bottlenecks adopt the same CNN blocks as used in Conformer, and ViT follows the design of Transformer branch. Models are grouped with comparable Params and MACs: Conformer-9, ResNet-152 and ViT-12; Conformer-6, ResNet-101 and ViT-9.
@@ 4.2.1 Benchmark evaluation
4.2.2 Enhanced prediction on conformational epitopes
To evaluate the ability of each method in predicting linear and conformational epitopes, we computed the F1 scores separately on the test set. As shown in Table 2, BConformeR achieves an F1 score of 0.105 on discontinuous epitopes, significantly higher than baseline methods that typically report these scores below 0.050. This illustrates that BConformeR is more effective at identifying non-contiguous epitope fragments.
@@ 4.3.1 Overview
To further assess the contribution of each module to linear and conformational epitope prediction, we ablated the dual-branch Conformer into standalone ResNet and ViT variants. We first reduced the depth of Conformer-12 (BConformeR) to evaluate the effectiveness of the original architecture. Then, ResNet-152 and ResNet-101, as well as ViT-12 and ViT-9, were selected based on comparable parameters (Params) and multiply-accumulate operations (MACs) to Conformer-9 and Conformer-6 respectively. The F1 score was used as the primary evaluation metric for both linear and discontinuous epitope prediction.
@@ 4.3.2 Depth reduction
To evaluate the impact of model depth, we progressively reduced the number of layers in the Conformer architecture. As shown in Table 3, Conformer-12 achieves the highest F1 scores on both linear (0.672) and discontinuous (0.110) epitopes. Reducing depth to Conformer-9 leads to a moderate drop in performance, with linear and discontinuous F1 scores decreasing by 4.0% (0.645) and 13.6% (0.095) respectively. Further reduction to Conformer-6 causes a more significant decline, especially in discontinuous epitope prediction, where the F1 score drops by 35.5% to 0.071. These demonstrate that our final model, BConformeR, achieves superior performance compared to shallower variants.
@@
Purely convolutional variants capture contiguous residue patterns efficiently and therefore achieve competitive scores on linear epitopes. However, their strictly local receptive field prevents them from modelling long-range residue couplings that dominate conformational epitopes. As shown in Table 3, ResNet-152 attains the highest linear F1 score (0.674) among all models, confirming its effectiveness on local patterns. However, its discontinuous F1 drops to 0.079, lower than Conformer-9 (0.095) and ViT-12 (0.086), indicating a weakness in capturing discontinuous dependencies.
By injecting a Transformer branch and exchanging information through FCUs, Conformer-9 achieves a discontinuous F1 score of 0.095, outperforming ResNet-152 (0.079), while maintaining a competitive linear F1 score of 0.645. This shows that integrating global context into convolutional layers can effectively improve discontinuous epitope prediction without sacrificing linear performance.
@@
Vision Transformer variants excel at aggregating sequence-wide information but lack explicit inductive bias for short-range motifs, often resulting in underfitting on linear stretches and noisy attention over densely packed regions. As shown in Table 3, ViT-9 achieves a discontinuous F1 of 0.082, surpassing ResNet-152 (0.079) and ResNet-101 (0.048). However, this comes at the cost of reduced performance on linear epitopes, where ViT-9 trails behind ResNet-101 (0.617 vs. 0.629).
Conformer mitigates this trade-off, as Conformer-9 outperforms ViT-12 on both linear (0.645 vs. 0.639) and discontinuous (0.095 vs. 0.086) F1 scores. Empirically, the hybrid design offers a more balanced solution for linear and discontinuous epitope prediction.

### v2

@@ Abstract
Accurate prediction of antibody-binding sites (epitopes) on antigens is crucial for vaccine design, immunodiagnostics, therapeutic antibody development, antibody engineering, research into autoimmune and allergic diseases, and for advancing our understanding of immune responses. Despite in silico methods that have been proposed to predict both linear (continuous) and conformational (discontinuous) epitopes, they consistently underperform in predicting conformational epitopes. In this work, we propose a conformer-based model trained on antigen sequences derived from 1,080 antigen-antibody complexes, leveraging convolutional neural networks (CNNs) to extract local features and Transformers to capture long-range dependencies within antigen sequences. Ablation studies demonstrate that CNN enhances the prediction of linear epitopes, and the Transformer module improves the prediction of conformational epitopes. Experimental results show that our model outperforms existing baselines in terms of PCC, ROC-AUC, PR-AUC, and F1 scores on both linear and conformational epitopes. The data and code are available at https://github.com/Jerry-alt19/BConformeR.
Keywords B-cell Epitope Prediction; Antigen–Antibody Interface; Conformer
@@ 1 Introduction
Figure 1: Overview of epitope classification and BConformeR prediction. (A) Epitopes, i.e., Antibody-binding sites, include conformational fragments (n=2,632) and linear regions (n=2,388). Data were derived from 1,432 antigen records (Appendix C.2). (B) Epitope prediction workflow. Antigen inputs are processed to predict linear and conformational epitopes, which enable applications in vaccine design and diagnostic development.
@@
Figure 2: BConformeR overview. (A) Preparation pipeline of the data. (Right) 1300 antigen-antibody FASTA files were processed using AlphaFold Multimer v3 to predict structures, with top-ranked structure for each was selected to extract epitopes. (Left) Antigen sequences were embedded by ESM-2, producing embeddings of dimension 1024×1280. (B) Conformer architecture. Stem module applies convolution and max-pooling to the original ESM-2 embeddings; The Conv block and Trans block are based on ResNet bottleneck and Vision Transformer modules respectively; Feature coupling units (FCU-Up and FCU-Down) connect the Conv and Trans blocks, with 11 FCU layers in total; The classifier linearly fuses features from both Conv and Trans branches.
@@
Here, we propose BConformeR, a Conformer-based model for continuous and discontinuous BCR epitope prediction. Antigen sequences from 1,180 Ag-Ab complexes were embedded using ESM-2, with 1,080 used for training and 100 for evaluation. In a blind test of 24 antigen records, BConformeR outperforms existing baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including PCC, ROC-AUC, PR-AUC, and F1 score on both linear and conformational epitopes. To further dissect the model’s design, we ablated ResNet bottlenecks and ViT modules from Conformer, showing that convolutional layers improve linear epitope prediction, and Transformer modules are critical for capturing discontinuous epitope patterns.
@@ 2.1 Antibody-agnostic prediction models
Numerous models, trained on antigen–epitope pairs, have been proposed and refined to enhance epitope prediction (Appendix A). SEPPA-3.0 was the epitope prediction tool specifically designed for glycoproteins, based on the logistic regression classifier SEPPA3.0 . BepiPred-3.0 used embeddings from pretrained protein language models ESM-2 as input to a feedforward neural network (FFNN) to predict both linear and conformational B-cell epitopes BepiPred3.0 . Epitope3D modeled antigen surfaces as residue-level graphs and extracted distance-based graph signatures to train classifiers for conformational B-cell epitope prediction, with AdaBoost selected as the backbone Epitope3D ; AdaBoost . ScanNet was an end-to-end geometric deep learning model that learned interpretable residue-level features directly from 3D structural neighborhoods ScanNet . SEMA-1D 2.0 was constructed by adding a fully-connected linear layer on an ensemble of five ESM-2 models, while SEMA-3D 2.0 was developed using an ensemble of five pre-trained Structure-aware Protein language models (SaProt) SEMA2.0 ; SaProt . DiscoTope-3.0 utilized inverse folding representations from ESM-IF1 and was trained on both predicted and solved structures using a positive-unlabelled ensemble strategy, enabling structure-based B-cell epitope prediction at scale DiscoTope3.0 ; ESM-IF1 .
@@ 3.1.2 Epitope classification
Generally, epitopes are separated into multiple Regions (R) and Fragments (F) for each antigen sequence. Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues. Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl (Appendix B). Fragments, which do not contain enough antibody-contacting residues to qualify as a region, are considered discontinuous in our study (Fig. 1A).
@@
Table 1: Architecture of BConformeR (BConformer-12), a 1D adaptation of Conformer Conformer21 . The MHSA-12 indicates multi-head self-attention with 12 heads operating on 1536-dimensional embeddings (128 dimensions per head). $\mathrm{Conv}_{\mathrm{cls}}$ denotes the 1D convolutional classifier operating on CNN features, and $\mathrm{Trans}_{\mathrm{cls}}$ denotes the linear classifier operating on Transformer outputs. The final prediction is obtained by a weighted fusion of the two.
@@ 3.2.2 CNN branch
The channel dimension is first downsampled, then expanded by a factor of 4 within each bottleneck. The first convolution block contains a single bottleneck, while subsequent blocks include two bottlenecks: one applied before the Transformer branch, and the other afterward. Since convolution kernels slide over representation sequences with overlap, the CNN blocks are well-suited to capture residue-level contextual information, providing local features for the Transformer branch.
@@ 3.2.3 Transformer branch
where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 12) heads Attention , and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$. The $\mathrm{DropPath}$ operator implements stochastic depth regularization.
@@ 3.3.1 AgIoU
To evaluate the overlap between the prediction and real epitopes, we use the Antigen-level Intersection over Union (AgIoU), which is defined as the ratio of the number of correctly predicted epitope residues to the total number of residues that are either labeled as epitope in the ground truth or prediction as such (Appendix D):
@@ 4.1.1 Dataset
We collected 1,300 Ag–Ab complex FASTA files with AlphaFold-2.3 colabfold version from AbEpiTope-1.0. Each complex was then modeled with AlphaFold2 Multimer v3, generating 30 structural predictions per complex and a total of 39,000 structures (Appendix C.1). For each complex, we selected the top-ranked model and excluded those with a pLDDT score below 0.82. Interface residues were extracted based on a distance cutoff of 4 Å between heavy atoms on the antigen and antibody chains (1). To enhance data quality and limit parameter overhead, we removed complexes containing insufficient epitope residues or antigens exceeding 1024 residues in length. Finally, the dataset contains 1,180 Ag-Ab complexes in total: 1,080 used for training BConformeR, and 100 set aside for evaluation and ablation. And the antigen chains annotated with epitope labels serve as the input.
@@
SEPPA-3.0 0.100 ± 0.020 0.158 ± 0.009 0.237 ± 0.025 0.058 ± 0.014 0.067 0.088 ± 0.005
BepiPred-3.0 0.220 ± 0.010 0.276 ± 0.007 0.209 ± 0.011 0.194 ± 0.003 0.215 0.155 ± 0.003
SEMA-1D 2.0 0.500 ± 0.030 0.270 ± 0.005 0.213 ± 0.006 0.188 ± 0.002 0.217 0.147 ± 0.002
DiscoTope-3.0 0.300 ± 0.020 0.303 ± 0.007 0.231 ± 0.004 0.227 ± 0.007 0.276 0.167 ± 0.002
BConformeR 0.330 ± 0.020 0.349 ± 0.004 0.262 ± 0.003 0.296 ± 0.003 0.347 0.203 ± 0.002
Table 2: Performance comparison of different methods on a blind test of 24 antigen records. F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively; Thresholds were selected following those of baseline methods (Appendix E.2); Standard deviation calculated over a range of thresholds.
Due to the differences in training and test datasets across existing epitope prediction methods, the selection of a blind test set was important for fair evaluation. Therefore, we also extracted 24 antigen records from the test sets of SEMA-1D 2.0 and DiscoTope-3.0, all of which are blind to BConformeR.
@@
BConformer-12 0.440 0.477 0.458 0.304 0.360 0.330 302.24 309.93
BConformer-9 0.428 0.463 0.445(-2.8%) 0.297 0.322 0.309(-6.4%) 230.27 236.12
BConformer-6 0.423 0.416 0.419(-8.5%) 0.249 0.296 0.270(-18.2%) 158.30 162.31
ResNet1D-152 0.442 0.484 0.462(+0.9%) 0.255 0.291 0.272(-17.6%) 242.44 248.62
ResNet1D-101 0.417 0.430 0.423(-7.6%) 0.242 0.238 0.240(-27.3%) 179.24 183.80
ViT1D-12 0.429 0.435 0.432(-5.7%) 0.269 0.334 0.298(-9.7%) 230.24 234.42
ViT1D-9 0.403 0.428 0.415(-9.4%) 0.276 0.311 0.292(-11.5%) 173.57 176.32
Table 3: Ablation study of different models evaluated on 100 antigen records (Appendix E.1). BConformer-12 corresponds to BConformeR, and BConformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to BConformer-12/9. ResNet bottlenecks adopt the same CNN blocks as used in Conformer, and ViT follows the design of Transformer branch (Appendix F). Models are grouped with comparable Params and MACs: BConformer-9, ResNet1D-152 and ViT1D-12; BConformer-6, ResNet1D-101 and ViT1D-9.
@@ 4.2.1 Benchmark evaluation
4.2.2 Enhanced prediction on continuous and discontinuous epitopes
To evaluate the ability of each method in predicting continuous and discontinuous epitopes, we computed the F1 scores separately on the test set. As shown in Table 2, BConformeR achieves F1 scores of 0.349 and 0.262 on linear and discontinuous epitopes, respectively, higher than all baseline methods. This illustrates that BConformeR is more effective at identifying both regions and fragments of epitopes.
@@ 4.3.1 Overview
To further assess the contribution of each module to linear and conformational epitope prediction, we ablated the dual-branch Conformer into standalone ResNet and ViT variants. We first reduced the depth of BConformer-12 (BConformeR) to evaluate the effectiveness of the original architecture. Then, ResNet1D-152 and ResNet1D-101, as well as ViT1D-12 and ViT1D-9, were selected based on comparable parameters (Params) and multiply-accumulate operations (MACs) to BConformer-9 and BConformer-6 respectively. The F1 score was used as the primary evaluation metric for both linear and discontinuous epitope prediction.
@@ 4.3.2 Depth reduction
To evaluate the impact of model depth, we progressively reduced the number of layers in the Conformer architecture. As shown in Table 3, reducing depth to BConformer-9 leads to a moderate drop in performance, with linear and discontinuous F1 scores decreasing by 2.8% (0.445) and 6.4% (0.309) respectively. Further reduction to BConformer-6 causes a more significant decline, especially in discontinuous epitope prediction, where the F1 score drops by 18.2% to 0.270. These demonstrate that our final model, BConformeR, achieves superior performance compared to shallower variants.
@@
Purely convolutional variants capture contiguous residue patterns efficiently and therefore achieve competitive scores on linear epitopes. However, their strictly local receptive field prevents them from modelling long-range residue couplings that dominate conformational epitopes. As shown in Table 3, ResNet1D-152 attains the highest linear F1 score (0.462) among all models, confirming its effectiveness on local patterns. However, its discontinuous F1 drops to 0.272, lower than BConformer-9 (0.309) and ViT1D-12 (0.298), pointing to difficulty in capturing discontinuous dependencies.
By injecting a Transformer branch and exchanging information through FCUs, BConformer-9 achieves a noticeably higher discontinuous F1 than ResNet1D-152, while keeping a linear F1 of 0.445 with only a minor drop. This shows that substituting part of the convolutional layers with a Transformer modules improves discontinuous epitope prediction with little sacrifice in linear performance.
@@
Vision Transformer variants excel at aggregating sequence-wide information but lack explicit inductive bias for short-range motifs, often resulting in underfitting on linear stretches and noisy attention over densely packed regions. As shown in Table 3, ViT1D-9 achieves a discontinuous F1 of 0.292, surpassing ResNet1D-101 (0.240), ResNet1D-152 (0.272) and BConformer-6 (0.270). However, this comes at the cost of reduced performance on linear epitopes, where ViT1D-9 (0.415) trails behind ResNet1D-101 (0.423) and BConformer-6 (0.419).
Conformer mitigates this trade-off, as BConformer-9 outperforms ViT1D-12 on both linear and discontinuous F1 scores. Empirically, the hybrid design offers a more balanced solution for linear and discontinuous epitope prediction.